OpenStudy (anonymous):

Greta has 45 minutes before her job interview. She is 15 miles away from the interview. How fast must she drive to be on time for the interview? (Hint: Convert the minutes to a fraction of an hour.)

OpenStudy (anonymous):

45 min is 3/4 of an hour so 15 miles in 45 min is: 15/(3/4)= 60/3=20 miles/hour